AO Commissioning

02:51 Continuing fallout from SAN failure.

04:01 Shells set and ready to go.

05:22 Heading to AO_53

05:37 Flaky Proximity sensor on LFBG Rotator

06:36 Sensor Fixed Ω

06:47 Collimating

06:51 Moving to M34 Luci Filters Brγ and K

07:19 Having Difficulty Homing SX Bayside Stages.

07:51 Loop Opened a few min ago. Seeing between 1.2 and 1.5.

08:27 Removed Actuator 223, it had very bad position measurements. I'll attach some images later.

08:45 Right side only again.

09:20 Working out position for the center of LUCI2.

09:47 Had to recalculate matrix to get GOPT working again.

09:55 GOPT functioning, but ratio is really low ~0.14-0.19.

10:00

Loop opened during offset because an edge rolled. This has been the 2nd time since GOPT was fixed, once with the regular reconstructor table and once in "Bad Seeing Mode". All failures were due to the loop not staying closed. There have been something like 10 so far.

I set the TTM to exactly 989.61 Hz.

10:11 Succsessful preset GOPT ratio 0.22-0.3

10:23 Off axis bin 2 star, I'm doing GOPT anyway. Ratio 0.08-0.12, seeing quite bad > 1.5. It was only a single iteration of DICE 5. We'll need to do some more.

10:34 We had a failure, due to high force or a bad actuator # 560. I wasn't watching the GUI.

The AdSec Control Gui has crashed several times today. It crashes when the shell is set, or changes state.

10:37 Repeting Off axis, bin 2 star. Doing 10" DICE 5 X 4 iterations. Preset magnitude was 10.05. WFS reads 10.9.

This camera has enough NCPA that I think GOPT might be worth while, even in bin 2.

10:52 Failed again, possibly due to a bad actuator # 560. We are going to move to the potential exoplanet object for Al.

Bin 4, Measured Magnitude 14.6, 311.87 Hz.

Preset Failed, probably took too long and timed out.

11:26 Successful preset, after reducing the number of iterations in Optimize Gain.

We will use the bright source itself as the reference star. It's 8th mag.

11:37 We are starting a new preset.

989.61 Hz. 8.9 Magnitude. Loop opened with Good seeing table. We've been having trouble with this all night. We are going back to the poor seeing table.

11:56 Actuator 560 Jumped for the 3 time tonight. i took it out.

12:16 Successfully started a preset. Bin 1 989 Hz, 8.9 Mag. Seeing between 0.9-1.1.

12:20 We had a hiccup in Barry's script, due to a problem with LUCI, but we were able to resume without having to do a new preset.

12:34The seeing suddenly jumped from 0.9 to 1.3 and the loop couldn't keep up.

12:36 Resumed

15:56 Finished script, ready to close-up, after Steve does a quick test.

-- %USERSIG{GregoryTaylor - 2016-10-06}%
I Attachment Action Size Date Who Comment
IMG_0572.JPGJPG IMG_0572.JPG manage 995 K 06 Oct 2016 - 23:40 GregoryTaylor Bad Pattern in Current
IMG_0573.JPGJPG IMG_0573.JPG manage 1 MB 06 Oct 2016 - 23:41 GregoryTaylor Actuator 223
IMG_0574.JPGJPG IMG_0574.JPG manage 730 K 06 Oct 2016 - 23:42 GregoryTaylor Shack Hartmann Pattern
Topic revision: r3 - 06 Oct 2016, GregoryTaylor
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback